OpenStudy (anonymous):

HELP ME PLEASE. Donna's company reimburses her expenses on food, lodging, and conveyance during business trips. The company pays $50 a day for food and lodging and $0.60 for each mile traveled. Donna drove 200 miles and was reimbursed $1620. Part A: Create an equation that will determine the number of days x on the trip (3 points) Part B: Solve this equation justifying each step with an algebraic property of equality. (6 points) Part C: How many days did Donna spend on this trip? (1 points)

OpenStudy (texaschic101):

(A) 50x + .60(200) = 1620 (B) 50x + .60(200) = 1620 <-- distributive property 50x + 120 = 1620 50x + 120 - 120 = 1620 - 120 <--subtraction property 50x = 1500 50x/50 = 1500/50 <--division property x = 30 (C) Donna spent 30 days on her trip.